Draw 10 times, WITH replacement, from a box of 100 tickets with 80 of them numbered 1 and the rest numbered 0, and add up the numbers on the tickets that are drawn. What is the

distribution?
Answer #1

Here , the probability of drawing a ticket numbered 1 is p= 80/100=0.8 and the ticket numbered 0 is =(1-p)= 0.2

The each trials are Bernoulli trials as all are independent with 1 as success and 0 as failure.

And probability of success=0.8

So, sum of the ten drawn i.e. sum of 10 Bernoulli trials follow binomial distribution with probability of success p=0.8 and n=10.